OpenStudy (anonymous):

Multiply: 5xy(2x - y)

OpenStudy (anonymous):

\[\left(5xy\cdot2x\right)-\left(5xy\cdot y\right)=\,\,?\]Your move.

OpenStudy (anonymous):

10x^2 y - 5x y^2 ??

OpenStudy (anonymous):

Yeah!